Solve the compound inequality Please.
I have been trying and I can not seem to figure out how to set the problem up to solve it.



(2x-1)/(4)<=-4 or (2x-1)/(4)>=9

Answer by Alan3354(69443)

(2x-1)/(4)<=-4
2x-1 <= -16
2x <= -15
x <= - 7.5
------------
or (2x-1)/(4)>=9
2x-1 >= 36
2x >= 37
x >= 18.5
-----------
x <= -7.5 or x >= 18.5